Suppose a firm operating in a perfectly competitive industry has costs in the short run given by:

SRTC = 8 + 1/2Q^2 and therefore MC q.



if the minimum point of the short-run ATC curve for all firms(existing and potential)is also the minimum point of the long run average cost curve (LRAC), calculate the long-run equilibrium price, market quantity, and firm quantity. What is the long-run equilibrium number of firms in the industry?


1
Expert's answer
2021-11-22T10:00:15-0500

Solution:

SRTC = 8 + 1/2Q2                         MC = q

Derive SRATC:

SRATC = "\frac{TC}{Q} = \frac{8 + 0.5Q^{2} }{Q} = \frac{8}{Q} + 0.5Q"


Derive the derivative of the average total cost:

"\frac{\partial SRATC} {\partial q} = \frac{-8} {q^{2} } + 0.5"

Set the derivative equal to zero and solve for q:

"\frac{-8} {q^{2} } + 0.5 = 0"

0.5q2 = 8

q2 = 16

q = 4


Long run equilibrium price: AC = "\frac{8} {q } + 0.5"

AC = "\frac{8} {4 } + 0.5(4)" = 2 + 2 = 4

Long run equilibrium price = 4

SRTC = 8 + 1/2Q2 = 8 + 0.5(4)2 = 8 + 8 = 16

The long-run equilibrium number of firms in the industry = "\frac{16} {4}" = 4 Firms